A rubber pad acts as an elastic spring.when a mass of 150g is placed on it a compression of 1.5cm is produced.
The spring mass system begins to oscillate when the mass is gently tapped downwards.
Calculate the frequency of oscillation.
1
Expert's answer
2011-03-17T07:16:04-0400
The period of the system: T = 2π (m/k)1/2
k = F/x = 0.15 [kg]* 10[N/kg]/0.015[m] = 100 [N/m]
f = 1/T = 1/2π (k/m)1/2 = 4.1 s-1.